# Stopping Rules for Stochastic Gradient Descent via Anytime-Valid Confidence Sequences

### v1

@@ 1.3 Related Work
Stochastic approximation and asymptotic normality. Stochastic approximation provides the theoretical foundation for stochastic gradient methods. The seminal paper [robbins1951stochastic] established almost sure convergence of iterative schemes under diminishing stepsizes, while [kushner2003stochastic] developed a comprehensive asymptotic theory based on ordinary differential equations, covering stability, asymptotic normality, and efficiency. A key advance showed that appropriately averaged iterates achieve asymptotically optimal covariance, providing a principled justification for studying weighted averages (which we adopt in this paper) of SGD trajectories [polyak1992acceleration]. This asymptotic perspective was sharpened in [duchi2021asymptotic], which characterized minimax and asymptotically efficient rates for stochastic optimization, including constrained formulations, which was further extended to nonsmooth stochastic approximation in [davis2024asymptotic]. Together, these results motivate the use of averaged iterates and enable asymptotic confidence regions for optimal solutions based on limiting normality. However, such guarantees are inherently asymptotic: they do not yield finite-time confidence sets, nor do they provide time-uniform validity under continuous monitoring or adaptive stopping.
@@
Fixed-horizon high-probability guarantees for SGD. A vast literature develops nonasymptotic convergence guarantees for stochastic gradient (and mirror-descent-type) methods that hold with high probability at a pre-specified iteration horizon. Early robust frameworks for stochastic convex optimization established fixed-confidence bounds under bounded or sub-Gaussian noise assumptions, providing optimal oracle complexity guarantees for stochastic approximation and stochastic programming [nemirovski2009robust, lan2012optimal]. Parallel developments in machine learning emphasized refined nonasymptotic analyses of stochastic approximation algorithms, including explicit bias-variance tradeoffs and finite-time rates for averaged iterates [moulines2011non, bach2013non, dieuleveut2016nonparametric]. Subsequent work sharpened fixed-horizon high-probability bounds across a range of regimes, including nonsmooth and strongly convex objectives, and clarified when optimal rates can be achieved by averaging versus last-iterate solutions [rakhlin2012making, shamir2013stochastic, harvey2019tight, harvey2019simple, jain2021making]. Complementary information-theoretic lower bounds characterize the fundamental limits of stochastic convex optimization at fixed confidence levels [agarwal2012information]. More recent contributions extend fixed-horizon high-probability guarantees beyond light-tailed noise models, developing robust stochastic gradient methods under heavy-tailed gradients via techniques such as clipping, truncation, and median-of-means constructions [gorbunov2020stochastic, cutkosky2021high, davis2021low]. Together, these works provide sharp finite-time guarantees at prescribed horizons and confidence levels. However, these guarantees are intrinsically fixed-time: they do not yield confidence sets or optimality certificates that remain valid uniformly over time under continuous monitoring or adaptive stopping without additional union bounds or restarts, motivating the time-uniform approach pursued in this paper.
@@
Time-uniform inference and confidence sequences. Time-uniform (also called anytime-valid) inference provides uncertainty bounds that remain valid simultaneously over all times and are therefore robust to continuous monitoring and optional stopping. A foundational tool is Ville’s inequality for nonnegative supermartingales [ville1939], which converts supermartingale constructions into time-uniform deviation guarantees. Early statistical instances of this idea include classical confidence sequences for basic parameters [darling1967confidence], as well as boundary-crossing results for martingales [robbins1970boundary]. Modern developments systematize these principles through mixture and stitching arguments, yielding sharp time-uniform concentration inequalities that adapt to variance-type processes [howard2020time, howard2021time]; see also the broader literature on self-normalized processes [de2009self]. More recently, these ideas have been reframed through e-values and game-theoretic statistics, providing a unified perspective on optional-stopping-safe inference and a comprehensive entry point to the literature [ramdas2023game, ramdas2025hypothesis]. While this framework has been developed primarily for sequential testing and estimation, it has begun to influence learning and decision-making problems, including anytime-valid inference for bandits and betting-based estimation [waudby2024anytime, waudby2024estimating]. Our work leverages this time-uniform machinery in an optimization context, using supermartingale mixtures to construct suboptimality certificates that remain valid under arbitrary, data-dependent stopping.
@@
Closest related work. The works most closely related to ours study anytime or time-uniform guarantees for stochastic approximation, but differ substantially in assumptions and in the form of the resulting guarantees. Classical results such as [rakhlin2012making] establish sharp high-probability convergence rates for SGD under strong convexity, bounded stochastic gradients, and fixed stepsizes, with guarantees stated for a pre-specified iteration horizon. As such, they do not support data-dependent stopping times. More recent work [feng2023anytime] derives anytime high-probability bounds for stochastic gradient methods with momentum under convexity and smoothness assumptions, using algorithm-specific dynamics and stepsize choices motivated by continuous-time ODE analysis. While these results control the optimization error uniformly over time, they do not provide explicit constants or trajectory-observable confidence sequences, and therefore cannot be used to derive certified stopping rules. In contrast to both works, our results apply to plain projected SGD under general convexity, yield fully explicit and trajectory-dependent confidence sequences for suboptimality, and lead to provably correct stopping rules that remain valid under arbitrary adaptive monitoring.
@@ 2.2 Clarifying remarks and intuition
The mixture construction in Lemma 2.7 reconciles these goals by fixing a countable family of parameters a priori and combining the corresponding exponential supermartingales. This “method of mixtures” (also called pseudo-maximization) goes back to Robbins and Siegmund’s classical work on sequential boundary crossing and is developed systematically in the theory of self-normalized processes (see [de2009self]). Modern confidence-sequence constructions, such as the sub-$\psi$ mixture framework in [howard2020time, howard2021time] and the e-process approach in [waudby2024estimating], rely crucially on this principle. $\clubsuit$
@@
In short, the grid is chosen so that: (i) it spans the relevant scales of $\Sigma_{t}^{2}$ via exponentially spaced search points, and (ii) at least one $\lambda_{k}$ is guaranteed to be within a constant factor of the ideal minimizer, thereby achieving near-optimal behavior across all times $t$. This type of “scale-matching” intuition for geometric grids is common in mixture and stitching arguments for martingale concentration; see, for example, the heuristic discussion preceding Theorem 1 in [howard2021time] and the grid-selection heuristics in the sub-$\psi$ mixture framework in [howard2020time]. $\clubsuit$
@@
A recurring feature in anytime martingale bounds is the presence of iterated logarithm terms such as $\log\log t$ or $\log\log d$ (see, e.g., Robbins–Siegmund iterated LIL bounds and mixture inequalities in [howard2020time]). At first glance, one might therefore expect a confidence sequence for stochastic optimization to include a term of the form $\sqrt{\Sigma_{t}^{2}\log\log t}$ rather than the $\sqrt{\Sigma_{t}^{2}\log\log(\mathrm{e}+\Sigma_{t}^{2})}$ term appearing in Theorem 2.2. The difference comes from the square-summable stepsize assumption in Assumption 1.2 and the resulting definition of $\Sigma_{t}^{2}=\sum_{s=1}^{t}\sigma^{2}\eta_{s}^{2}Z_{s}$. In most sequential testing settings, $\Sigma_{t}^{2}$ grows linearly with $t$, so controlling $k$ in Step (iv) of Lemma 2.7 ultimately leads to iterated-logarithm terms in $\log t$. Here, however, $\{\eta_{s}\}_{s\geq 1}$ is square summable, so $\Sigma_{t}^{2}$ is bounded and converges to
@@
This phenomenon is well known in martingale boundary literature: confidence sequences expressed in terms of data-dependent variance proxies (predictable quadratic variation, conditional variance processes, or self-normalized sums) shrink faster than those written in terms of time alone; see [de2009self] and the “variance process versus time” discussion in [howard2021time, Section 3]. $\clubsuit$
@@ 3.1 Stopping-time complexity
The polynomial regime $\gamma\in(1/2,1)$ produces certified complexity $\tau_{\varepsilon}=O\Big(\varepsilon^{-1/(1-\gamma)}\Big)$, which matches the classical stochastic approximation tradeoff between stepsize decay and variance control. The harmonic schedule $\eta_{t}=\eta_{0}/t$, although valid for confidence-sequence stopping, leads to exponential complexity in $1/\varepsilon$ and is therefore unattractive in practice. Combined with Theorem 3.1, these results show that the stepsize schedule entirely governs the certified stopping rate, and that the confidence-sequence approach preserves standard stochastic approximation complexity guarantees while remaining valid under arbitrary, data-dependent optional stopping. $\clubsuit$

### v2

@@ 1.3 Related Work
Stochastic approximation and asymptotic normality. Stochastic approximation provides the theoretical foundation for stochastic gradient methods. The seminal paper [25] established almost sure convergence of iterative schemes under diminishing stepsizes, while [17] developed a comprehensive asymptotic theory based on ordinary differential equations, covering stability, asymptotic normality, and efficiency. A key advance showed that appropriately averaged iterates achieve asymptotically optimal covariance, providing a principled justification for studying weighted averages (which we adopt in this paper) of SGD trajectories [21]. This asymptotic perspective was sharpened in [9], which characterized minimax and asymptotically efficient rates for stochastic optimization, including constrained formulations, which was further extended to nonsmooth stochastic approximation in [6]. Together, these results motivate the use of averaged iterates and enable asymptotic confidence regions for optimal solutions based on limiting normality. However, such guarantees are inherently asymptotic: they do not yield finite-time confidence sets, nor do they provide time-uniform validity under continuous monitoring or adaptive stopping.
@@
Fixed-horizon high-probability guarantees for SGD. A vast literature develops nonasymptotic convergence guarantees for stochastic gradient (and mirror-descent-type) methods that hold with high probability at a pre-specified iteration horizon. Early robust frameworks for stochastic convex optimization established high-probability bounds under bounded or sub-Gaussian noise assumptions, providing optimal oracle complexity guarantees for stochastic approximation and stochastic programming [20, 18]. Parallel developments in machine learning emphasized refined nonasymptotic analyses of stochastic approximation algorithms, including explicit bias–variance tradeoffs and finite-time rates for averaged iterates [19, 2, 8]. Subsequent work sharpened fixed-horizon high-probability bounds across a range of regimes, including nonsmooth and strongly convex objectives, and clarified when optimal rates can be achieved by averaging versus last-iterate solutions [22, 27, 12, 13, 16]. Complementary information-theoretic lower bounds characterize the fundamental limits of stochastic convex optimization at fixed confidence levels [1]. More recent contributions extend fixed-horizon high-probability guarantees beyond light-tailed noise models, including heavy-tailed settings handled via gradient clipping and related robustification techniques [11, 3], as well as general boosting and robust-estimation wrappers that convert low-probability guarantees into high-confidence ones [5]. Together, these works provide sharp finite-time guarantees at prescribed horizons and confidence levels. However, these guarantees are intrinsically fixed-time: they do not yield confidence sets or optimality certificates that remain valid uniformly over time under continuous monitoring or adaptive stopping without additional union bounds or restarts, motivating the time-uniform approach pursued in this paper.
@@
Time-uniform inference and confidence sequences. Time-uniform (also called anytime-valid) inference provides uncertainty bounds that remain valid simultaneously over all times and are therefore robust to continuous monitoring and optional stopping. A foundational tool is Ville’s inequality for nonnegative supermartingales [28], which converts supermartingale constructions into time-uniform deviation guarantees. Early statistical instances of this idea include classical confidence sequences for basic parameters [4], as well as boundary-crossing results for martingales [26]. Modern developments systematize these principles through nonnegative supermartingale constructions and, in particular, mixture and stitching arguments, yielding sharp time-uniform concentration inequalities that adapt to variance-type processes [15]; see also related time-uniform Chernoff bounds based on exponential supermartingales [14] and the broader literature on self-normalized processes [7]. More recently, these ideas have been reframed through e-values and game-theoretic statistics, providing a unified perspective on optional-stopping-safe inference and a comprehensive entry point to the literature [24, 23]. While this framework has been developed primarily for sequential testing and estimation, it has begun to influence learning and decision-making problems, including anytime-valid inference for bandits and betting-based estimation [30, 29]. Our work leverages this time-uniform machinery in an optimization context, using supermartingale mixtures to construct suboptimality certificates that remain valid under arbitrary, data-dependent stopping.
@@
Closest related work. The works most closely related to ours study anytime or time-uniform guarantees for stochastic approximation, but differ substantially in assumptions and in the form of the resulting guarantees. Classical results such as [22] establish sharp high-probability convergence rates for SGD under strong convexity, bounded stochastic gradients, and fixed stepsizes, with guarantees stated for a pre-specified iteration horizon. As such, they do not support data-dependent stopping times. More recent work [10] derives anytime high-probability bounds for stochastic gradient methods with momentum under convexity and smoothness assumptions, using algorithm-specific dynamics and stepsize choices motivated by continuous-time ODE analysis. While these results control the optimization error uniformly over time, they do not provide explicit constants or trajectory-observable confidence sequences, and therefore cannot be used to derive certified stopping rules. In contrast to both works, our results apply to plain projected SGD under general convexity, yield fully explicit and trajectory-dependent confidence sequences for suboptimality, and lead to provably correct stopping rules that remain valid under arbitrary adaptive monitoring.
@@ 2.2 Clarifying remarks and intuition
The mixture construction in Lemma 2.7 reconciles these goals by fixing a countable family of parameters a priori and combining the corresponding exponential supermartingales. This “method of mixtures” (also called pseudo-maximization) goes back to Robbins and Siegmund’s classical work on sequential boundary crossing and is developed systematically in the theory of self-normalized processes (see [7]). Modern confidence-sequence constructions build on this principle in various forms, including explicit mixture and stitching frameworks such as the sub-$\psi$ approach in [15], as well as betting- and e-process-based methods that aggregate families of nonnegative supermartingales; see, [29]. $\clubsuit$
@@
In short, the grid is chosen so that: (i) it spans the relevant scales of $\Sigma_{t}^{2}$ via exponentially spaced search points, and (ii) at least one $\lambda_{k}$ is guaranteed to be within a constant factor of the ideal minimizer, thereby achieving near-optimal behavior across all times $t$. This “scale-matching” intuition for geometric discretizations is common in mixture and stitching arguments for martingale concentration; see, for example, the heuristic discussion surrounding Theorem 1 in [15] and the broader treatment of pseudo-maximization via mixtures in self-normalized processes [7]. $\clubsuit$
@@
A recurring feature in time-uniform martingale concentration is the appearance of iterated logarithm corrections, such as $\log\log t$; see, for example, classical boundary-crossing analyses of Robbins and Siegmund [26] and modern confidence-sequence treatments such as [15]. At first glance, one might therefore expect a confidence sequence for stochastic optimization to include a term of the form $\sqrt{\Sigma_{t}^{2}\log\log t}$ rather than the $\sqrt{\Sigma_{t}^{2}\log\log(\mathrm{e}+\Sigma_{t}^{2})}$ term appearing in Theorem 2.2. The difference comes from the square-summable stepsize assumption in Assumption 1.2 and the resulting definition of $\Sigma_{t}^{2}=\sum_{s=1}^{t}\sigma^{2}\eta_{s}^{2}Z_{s}$. In most sequential testing settings, $\Sigma_{t}^{2}$ grows linearly with $t$, so controlling $k$ in Step (iv) of Lemma 2.7 ultimately leads to iterated-logarithm terms in $\log t$. Here, however, $\{\eta_{s}\}_{s\geq 1}$ is square summable, so $\Sigma_{t}^{2}$ is bounded and converges to
@@
This phenomenon is well known in the martingale boundary literature: confidence sequences expressed in terms of data-dependent variance proxies (such as predictable quadratic variation, conditional variance processes, or self-normalized sums) adapt to the realized variability of the process and can be substantially tighter when the variance grows slowly or remains bounded, compared to bounds written purely in terms of time. See, for example, [7] and the discussion in [15, Section 3] on expressing time-uniform bounds in terms of intrinsic variance processes rather than the time index. $\clubsuit$
@@ 3.1 Stopping-time complexity
The polynomial regime $\gamma\in(1/2,1)$ produces certified complexity $\tau_{\varepsilon}=O\Big(\varepsilon^{-1/(1-\gamma)}\Big)$, which matches the classical stochastic approximation tradeoff between stepsize decay and variance control. The harmonic schedule $\eta_{t}=\eta_{0}/t$, although valid for confidence-sequence stopping, leads to exponential complexity in $1/\varepsilon$ and can therefore be unattractive in practice. Combined with Theorem 3.1, these results show that the stepsize schedule entirely governs the certified stopping-time rate (up to problem-dependent constants), and that the confidence-sequence approach preserves standard stochastic approximation complexity guarantees while remaining valid under arbitrary, data-dependent optional stopping. $\clubsuit$